A 50lb Boxer needs 3- IU/Kg of U-100 insulin twice daily. How much will the owner need to use over 4 weeks?
Answer #1

ANSWER.

* Calculation of amount of insulin use over 4 weeks.

           Ordered dose = 3 IU/kg twice daily

      Twice daily is two times a day

         So

             Ordered dose = 3 * 2

                                    = 6 IU/kg per day.

             Weight of the patient = 50 lb

    Convert lb to kg for calculation

            1 lb = 0.454 kg

     So

          Weight of the patient = 50 * 0.454

                                            = 22.7 kg.

     Dose required for the patient per day = 6 * 22.7

                                                                 = 136.2 IU per day.

      Dose required over 4 weeks = 136.2 * 28 (4 week)

                                                     = 3813.6 IU.